# ORFuzz: Fuzzing the "Other Side" of LLM Safety – Testing Over-Refusal

### v1

@@ Abstract
Large Language Models (LLMs) increasingly exhibit over-refusal—erroneously rejecting benign queries due to overly conservative safety measures—a critical functional flaw that undermines their reliability and usability. Current methods for testing this behavior are demonstrably inadequate, suffering from flawed benchmarks and limited test generation capabilities, as highlighted by our empirical user study. To the best of our knowledge, this paper introduces the first evolutionary testing framework, ORFuzz, for the systematic detection and analysis of LLM over-refusals. ORFuzz uniquely integrates three core components: (1) safety category-aware seed selection for comprehensive test coverage, (2) adaptive mutator optimization using reasoning LLMs to generate effective test cases, and (3) OR-Judge, a human-aligned judge model validated to accurately reflect user perception of toxicity and refusal. Our extensive evaluations demonstrate that ORFuzz generates diverse, validated over-refusal instances at a rate (6.98% average) more than double that of leading baselines, effectively uncovering vulnerabilities. Furthermore, ORFuzz’s outputs form the basis of ORFuzzSet, a new benchmark of 1,855 highly transferable test cases that achieves a superior 63.56% average over-refusal rate across 10 diverse LLMs, significantly outperforming existing datasets. ORFuzz and ORFuzzSet provide a robust automated testing framework and a valuable community resource, paving the way for developing more reliable and trustworthy LLM-based software systems. The code of this paper is available at: https://github.com/HotBento/ORFuzz.
@@ I Introduction
These components operate in an iterative fuzzing loop, enabling ORFuzz to automatically generate diverse and high-quality test cases that effectively expose over-refusal vulnerabilities across various safety categories. Our comprehensive evaluations demonstrate that ORFuzz significantly outperforms existing methods, achieving a validated over-refusal generation rate of 6.98%—more than double that of comparable baselines. We also find that various LLMs, owing to variations in training data and safety mechanisms, exhibit distinct over-refusal rates across safety categories. Furthermore, the application of ORFuzz has yielded ORFuzzSet, a new benchmark of 1,855 highly effective and transferable test cases, which itself demonstrates superior performance in triggering over-refusals across a wide array of LLMs.
@@
  • We construct and release a benchmark dataset of 1,855 high-quality test cases for over-refusal, generated and validated by ORFuzz. This dataset demonstrates high efficacy, triggering an average over-refusal rate of 63.56% across 10 diverse LLMs, serving as a robust resource for future testing and evaluation of LLMs.
@@ V-C Transferability Analysis and Benchmark Construction (RQ4)
Figure 4: Left: ORR of new benchmark ORFuzzSet vs. existing datasets on 10 LLMs. Right: t-SNE of ORFuzzSet’s semantic diversity.
@@
Leveraging these findings, we curated a new benchmark dataset, termed ORFuzzSet, comprising 1,855 queries from ORFuzz’s output that successfully triggered over-refusal in at least three of the five initial target LLMs. We then evaluated ORFuzzSet on an expanded set of 10 diverse LLMs (the original five plus Qwen2.5-3B-Instruct, Qwen2.5-14B-Instruct (re-evaluated), Deepseek-R1-Distill-Qwen-7B, Deepseek-R1-Distill-Llama-8B, and Deepseek-R1-Distill-Qwen-14B). For comparison, we also evaluated existing datasets (XSTest, OR-Bench, COR) on these 10 LLMs. As shown in Fig. 4 (left), ORFuzzSet consistently achieves the highest average ORR (63.56%) across all 10 models, significantly outperforming prior benchmarks and validating its broad effectiveness. Interestingly, within the Qwen series, over-refusal propensity (on ORFuzzSet) did not strictly correlate with model size: Qwen2.5-7B (70.94% ORR) was comparable to Qwen2.5-14B (also 70.94% ORR) and notably higher than Qwen2.5-3B (51.10% ORR). The t-SNE visualization in Fig. 4 (right) illustrates the extensive semantic distribution of samples in ORFuzzSet, indicating a high degree of diversity. We also count the number of samples in each safety category, as shown in Table V. The results show that ORFuzzSet covers all 8 safety categories. Among them, the most common categories are Crimes and Illegal Activities (CI) (23.18%) and Physical and Mental Health (PM) (18.71%), indicating that over-refusal samples in these categories are more likely to be transferable across models.
Answer to RQ4 ORFuzz can generate over-refusal samples with notable transferability. The curated benchmark dataset, ORFuzzSet (1,855 samples), derived from these transferable instances, demonstrates superior effectiveness (average 63.56% ORR on 10 LLMs) and high diversity, offering a valuable new resource for robustly evaluating LLM over-refusal.
@@ VI Conclusion
This paper presents ORFuzz, the first testing framework addressing LLM over-refusal and current detection limitations to the best of our knowledge. ORFuzz combines category-aware seed selection, adaptive mutator optimization, and OR-Judge, a human-aligned judge model whose accurate perception of input toxicity and refusal our evaluations confirm. Experiments show ORFuzz significantly outperforms baselines in generating diverse over-refusal samples (6.98% average ORR), with all components proving crucial. ORFuzz’s transferable outputs form ORFuzzSet, a new 1,855-sample benchmark that demonstrates superior effectiveness (63.56% average ORR across 10 LLMs) over prior datasets. ORFuzz and ORFuzzSet provide a robust automated testing framework and a key resource for enhancing LLM dependability through rigorous over-refusal assessment.

### v2

@@ Abstract
Large Language Models (LLMs) have been found to show over-refusal problems—erroneously rejecting benign queries due to overly conservative safety measures—a critical functional flaw that undermines their reliability and usability. Current methods for testing this behavior are demonstrably inadequate, suffering from flawed benchmarks and limited test generation capabilities, as highlighted by our empirical user study. To the best of our knowledge, this paper introduces the first evolutionary testing framework, ORFuzz, for the systematic detection and analysis of LLM over-refusals. ORFuzz uniquely integrates three core components: (1) safety category-aware seed selection for comprehensive test coverage, (2) adaptive mutator optimization using reasoning LLMs to generate effective test cases, and (3) OR-Judge, a human-aligned judge model validated to accurately reflect user perception of toxicity and refusal. Our extensive evaluations demonstrate that ORFuzz generates diverse, validated over-refusal instances at a rate (6.98% average) more than double that of leading baselines, effectively uncovering vulnerabilities. Furthermore, ORFuzz’s outputs form the basis of ORFuzzSet, a new benchmark of 1,786 highly transferable test cases that achieves a superior 57.37% average over-refusal rate across 14 diverse LLMs, significantly outperforming existing datasets. ORFuzz and ORFuzzSet provide a robust automated testing framework and a valuable community resource, paving the way for developing more reliable and trustworthy LLM-based software systems. The code of this paper is available at: https://github.com/HotBento/ORFuzz.
@@ I Introduction
These components operate in an iterative fuzzing loop, enabling ORFuzz to automatically generate diverse and high-quality test cases that effectively expose over-refusal vulnerabilities across various safety categories. Our comprehensive evaluations demonstrate that ORFuzz significantly outperforms existing methods, achieving a validated over-refusal generation rate of 6.98%—more than double that of comparable baselines. We also find that various LLMs, owing to variations in training data and safety mechanisms, exhibit distinct over-refusal rates across safety categories. Furthermore, the application of ORFuzz has yielded ORFuzzSet, a new benchmark of 1,786 highly effective and transferable test cases, which itself demonstrates superior performance in triggering over-refusals across a wide array of LLMs.
@@
  • We construct and release a benchmark dataset of 1,786 high-quality test cases for over-refusal, generated and validated by ORFuzz. This dataset demonstrates high efficacy, triggering an average over-refusal rate of 57.37% across 14 diverse LLMs, serving as a robust resource for future testing and evaluation of LLMs.
@@ V-C Transferability Analysis and Benchmark Construction (RQ4)
Figure 4: Left: ORR of new benchmark ORFuzzSet vs. existing datasets on 14 LLMs. Right: t-SNE of ORFuzzSet’s semantic diversity.
@@
Leveraging these findings, we curated a new benchmark dataset, termed ORFuzzSet, comprising 1,786 queries from ORFuzz’s output that successfully triggered over-refusal in at least three of the five initial target LLMs. We then evaluated ORFuzzSet on an expanded set of 14 diverse LLMs (the original five plus Qwen2.5-3B-Instruct, Qwen2.5-14B-Instruct (re-evaluated), Deepseek-R1-Distill-Qwen-7B, Deepseek-R1-Distill-Llama-8B, Deepseek-R1-Distill-Qwen-14B, GPT-4o, o3, o3-mini, and o4-mini). For comparison, we also evaluated existing datasets (XSTest, OR-Bench, COR) on these 14 LLMs. As shown in Fig. 4 (left), ORFuzzSet consistently achieves the highest average ORR (57.37%) across all 14 models, significantly outperforming prior benchmarks and validating its broad effectiveness. Interestingly, within the Qwen series, over-refusal propensity (on ORFuzzSet) did not strictly correlate with model size: Qwen2.5-7B (70.94% ORR) was comparable to Qwen2.5-14B (also 70.94% ORR) and notably higher than Qwen2.5-3B (51.10% ORR). Another interesting observation is that reasoning models (e.g., o3, o3-mini, o4-mini) exhibit lower ORR on ORFuzzSet compared to their base counterparts (GPT-4o), suggesting that enhanced reasoning capabilities may help mitigate over-refusal tendencies. The t-SNE visualization in Fig. 4 (right) illustrates the extensive semantic distribution of samples in ORFuzzSet, indicating a high degree of diversity. We also count the number of samples in each safety category, as shown in Table V. The results show that ORFuzzSet covers all 8 safety categories. Among them, the most common categories are Crimes and Illegal Activities (CI) (23.18%) and Physical and Mental Health (PM) (18.71%), indicating that over-refusal samples in these categories are more likely to be transferable across models.
Answer to RQ4 ORFuzz can generate over-refusal samples with notable transferability. The curated benchmark dataset, ORFuzzSet (1,786 samples), derived from these transferable instances, demonstrates superior effectiveness (average 57.37% ORR on 14 LLMs) and high diversity, offering a valuable new resource for robustly evaluating LLM over-refusal.
@@ VI Conclusion
This paper presents ORFuzz, the first testing framework addressing LLM over-refusal and current detection limitations to the best of our knowledge. ORFuzz combines category-aware seed selection, adaptive mutator optimization, and OR-Judge, a human-aligned judge model whose accurate perception of input toxicity and refusal our evaluations confirm. Experiments show ORFuzz significantly outperforms baselines in generating diverse over-refusal samples (6.98% average ORR), with all components proving crucial. ORFuzz’s transferable outputs form ORFuzzSet, a new 1,786-sample benchmark that demonstrates superior effectiveness (57.37% average ORR across 14 LLMs) over prior datasets. ORFuzz and ORFuzzSet provide a robust automated testing framework and a key resource for enhancing LLM dependability through rigorous over-refusal assessment. Generally, our work enables two key applications for mitigating over-refusal in LLMs. First, the ORFuzzSet benchmark can be used for targeted fine-tuning, helping developers reduce over-refusal rates by exposing models to diverse benign queries that previously triggered erroneous refusals. Second, the OR-Judge model can serve as a real-time refusal checker, providing a “second opinion” on whether a refusal is justified for a given prompt. We hope our work inspires further research into this important yet underexplored aspect of LLM safety and usability.
Acknowledgment
This work is supported by the State Key Laboratory of Industrial Control Technology, China (Grant No.ICT2024C01), and the Fundamental Research Funds for the Central Universities, China (Grant No.2025ZFJH02). Additionally, the authors wish to thank the anonymous reviewers for their helpful comments.